# Satellite Autonomous Clock Fault Monitoring with Inter-Satellite Ranges Using Euclidean Distance Matrices

### v2

@@ 4.1.1 Definitions
Based on the graph and observed ranges $r_{ij}$, we define a weighted graph $G=\langle V,E,W\rangle$. A realization (embedding) of $G$ is $d$-dimensional Euclidean distance space, $\mathbb{R}^{d}$, is defined as a function $\Phi$ that maps $V$ into $\mathbb{R}^{d}$, such that for each edge $e=(v_{i},v_{j})\in E$, $W(e)=r_{ij}$ (jian_beyond_2010). A weighted graph $G=\langle V,E,W\rangle$ is called d-embeddable when there is a realization (embedding) $\Phi:V\rightarrow\mathbb{R}^{d}$ that maps the edges to points ($\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$) in $\mathbb{R}^{d}$ space.
We define “fault satellite” as a satellite that experiences a clock phase jump, and corrupts at least one range measurement containing itself. A weighted graph is fault disprovable if and only if the embeddability of $G$ implies that it contains no fault satellites, and the non-embeddability of $G$ implies that it contains at least one fault satellite.
A graph is called rigid if it has no continuous deformation other than rotation, translation, and reflection while preserving the relative distance constraints between the vertices (laman_graphs_1970). Otherwise, it is called a flexible graph. A graph is called k-vertex redundantly rigid if it remains rigid after removing any $(k-1)$ vertices.
@@ 4.1.2 Embeddability of Graphs with Fault Satellites
Suppose there is at most $k-1$ fault satellite on a graph $G=\langle V,E,W\rangle$ that is $k$-vertex ($k\geq 2$) redundantly rigid in $\mathbb{R}^{d}$. If $G$ is $d$-embeddable, then $G$ contains no fault vertex (satellite) with probability 1.
@@
Given a weighted graph $G$, if $G$ is $k$-vertex redundantly rigid ($k\geq 2)$, then $G$ is fault disprovable at probability 1, if there are at most $k-1$ fault satellites.
@@
For an arbitrary graph $G$, if $G$ is not embeddable, then $G$ contains at least one fault satellite. Therefore, with Proposition 4.1, if $G$ is $k$-vertex redundantly rigid ($k\geq 2)$, then $G$ is fault disprovable at probability 1, if there are at most $k-1$ fault satellites. ∎
@@ 4.2 Properties of the Euclidean Distance Matrix
In the previous section, we proved that by using vertex redundantly rigid graphs, we can detect faults by analyzing their embeddability. Analyzing the embeddability of an arbitrary graph is known to be NP-hard  (Hendrickson1992), but if the graph is fully connected, we can evaluate it by analyzing the ranks of the GCEDM. Note that fully connected graphs of more than 5 nodes are 2-vertex redundantly rigid in $\mathbb{R}^{3}$.
@@ 4.2.1 Geometric Centered Euclidean Distance Matrix
where $\mathbf{1}\in\mathbb{R}^{n}$ is the ones vector. When no fault or noise is present $(m=0)$, the $\mathbf{G}^{n,d,0}$ will be positive semi-definite, and its rank will satisfy $\mathrm{rank}(\mathbf{G}^{n,d,0})=\mathrm{rank}(\mathbf{X}^{\top}\mathbf{X})\leq d$ (dokmanic2015euclidean).
@@ 4.2.2 Rank of the Euclidean Distance Matrices with Fault Satellites
Knowles2023 observed that there are more than three nonzero singular values when the GCEDM is constructed from an EDM with fault satellites. An example is provided in Figure 4. However, mathematical proofs were not provided in their paper. In this section, we provide and prove several properties of the EDM and GCEDM that are corrupted with faults and measurement noises.
@@
Figure 4: The log plot of the singular values of an example geometric centered EDM for $n=12,d=3$. The singular values are ordered in the descending order of its magnitude. The blue circles show the singular values when there is no fault or noise ($\mathbf{G}^{12,3,0}$), the orange squares show the singular values when there is noise (of scale $\sigma_{w}=10^{-6}$) added to the range measurements ($\tilde{\mathbf{G}}^{12,3,0}$), and the green crosses show the singular values when both noise and fault (of scale $\bar{b}=10^{-3}$) is added ($\tilde{\mathbf{G}}^{12,3,2}$). The 12 points are identical for all cases, which are sampled randomly inside a unit cube. When noise is present, the singular values increase except for the last singular value, which is close to zero (theoretically zero). When $m$ faults with magnitudes that are larger than the noise magnitude are added, the singular values up to the $d+2m$ index increase compared to the non-fault case.
@@ 4.2.3 Distribution of the 4th Singular Value of the Geometric Centered EDM
where $\lambda_{i}$ is the $i$th singular value of the geometric centered EDM. If the ISR measurements are completely noiseless, we can detect if a fault satellite exists within the graph by checking if $\gamma_{\text{test}}$ of the GCEDMs is not 0, since GCEDMs have rank 3 when no fault exists. However, when the ISR measurements are noisy, the 4th singular value increase regardless of faults, as shown in Figure 4 and Corollary 4.4.1. Therefore, to detect faults under the presence of noises, we need to set a threshold based on the singular value distributions of the noisy (but non-fault) GCEDMs to separate the fault from the noise.
For $n=5$, we are able to compute the distribution of the test statistic analytically. The distribution of the test statistic $\gamma_{\text{test}}/s$ follows a centralized chi-squared distribution with degree of freedom 1, as follows:
$\frac{\gamma_{\text{test}}}{s}$ $\sim\chi^{2}\left(1,0\right)\quad\text{for }n=5$ (46)
where the scale $s$ is given by the following equation:
$s$ $=2\sum_{i=1}^{5}\sum_{j=1}^{5}\left(\sigma_{m}\tilde{\mathbf{D}}_{ij}\right)^{2}\cdot\|\mathbf{\hat{U}}_{i,:}\|^{2}\cdot\|\mathbf{\hat{V}}_{j,:}\|^{2}\quad\quad\mathbf{\hat{U}}_{i,:},\mathbf{\hat{V}_{j,:}}:\text{$i$th and$j$th row of }\mathbf{\hat{U}},\mathbf{\hat{V}}\in\mathbb{R}^{5\times 2}$ (47)
@@
Above, $\tilde{\mathbf{G}},\tilde{\mathbf{D}}\in\mathbb{R}^{5times5}$ are the (noisy) GCEDM and the distance matrix constructed from the measured ranges, respectively. The derivation is given in Appendix F. The scale of the distribution is proportional to the noise magnitude $\sigma_{m}^{2}$ (Figure 5). Equation (47) can be modified to compute the scale when the variance of noise is different for each measurement, by replacing $\sigma_{m}$ with $(\sigma_{m})_{ij}$ (the noise for each range measurement).
When there is a clock jump at satellite $s_{f}$, the distribution of the test statistic $({\gamma_{\text{test}}})_{s_{f}}$ follows a non-central chi-squared distribution with 1 degree of freedom, with its scale $s$ given by the same equation as above, and non-centrality parameter $\lambda_{s_{f}}$:
$\frac{({\gamma_{\text{test}}})_{s_{f}}}{s}$ $\sim\chi^{2}\left(1,\lambda_{s_{f}}\right)\quad\text{for }n=5$ (51)
@@
$\lambda_{s_{f}}$ $=\frac{1}{s}\|\hat{\mathbf{U}}^{\top}\left(\mathbf{F}\circ\mathbf{D}\right)\hat{\mathbf{V}}\|_{F}^{2}$ (52)
$\mathbf{F}$ =$=\begin{cases}f_{ij}&i=s_{f}\ \text{or}\ j=s_{f}\\ 0&\text{otherwise}\end{cases}$ (53)
Therefore, when all biases added by faults have the same magnitude, the non-centrality parameter $\lambda$ is proportional to the square of the fault magnitude as shown in Figure 6 and Figure 7. The derivation is given in Appendix G.
$MDB(i)=\sqrt{\lambda_{i}\cdot s},\quad P[\chi^{2}(1,\lambda_{i})\leq\chi^{2}_{\alpha}(1,0)]=1-\gamma$ (54)
where $\alpha$ is the false alarm rate, $\gamma$ is the detection power, $s$ is the scale, provided in (47), and $\lambda_{i}$ is the non-centrality parameter for satellite $i$ fault case, provide in equation (52).
@@
Figure 5: The distribution of the test statistic $\gamma_{\text{test}}$ and its scale $s$ for different noise levels $\sigma_{m}$. The distribution of the test statistic $\gamma_{\text{test}}$ follows a chi-squared distribution with 1 degree of freedom. The scale $s$ is proportional to the noise magnitude $\sigma_{m}^{2}$.
@@ 4.3.1 Clique Listing
In order to construct an EDM and GCEDM, we need a set of range measurements between all pairs of $n$ satellites. However, due to the occultation by the planetary body and attitude constraints, the satellites in the constellation are not fully connected with ISRs in most cases. Therefore, the first step of the fault detection algorithm is to find a set of $k$-clique subgraphs, which are fully connected subgraphs of $k$ satellites. For the graph to be 2-vertex redundantly rigid, and to use the derived distritbution (46), we are interested in cliques with $k=5$. An example of 5-cliques subgraphs is shown in Figure 8.
Various exact algorithms to find all $k$-cliques have been proposed (Li2020OrderingHF). In this paper, we used the Chiba-Nishizeki Algorithm (Albo) due to its simplicity of implementation (Chiba1985). For each node $v$, Arbo expands the list of $k$-cliques by recursively creating a subgraph induced by $v$’s neighbors. The readers are referred to (Li2020OrderingHF) for the pseudocode and summary of the $k$-clique listing algorithm. Note that the clique finding algorithm does not have to be executed online. Since the availability of the ISRs can be predicted beforehand when scheduling the ISLs using a coarse predicted orbit, we can estimate the topology of the ISLs for future time epochs. We can compute the list of $k$-cliques of the predicted topologies for future time epochs (e.g., one orbit), and uplink them to satellites intermittently. If some of these links were actually not available in orbit for some reason, we can remove the $k$-cliques that contain the missing links from the list.
@@ 4.3.2 Online Fault Detection
$H_{i}:$ The satellite $\text{The satellite }i\text{ is a fault satellite}$
@@
    First, we compute the scaled test statistic $\gamma_{test,G^{\prime}}=\lambda_{4}^{2}/s$ for each 5-clique subgraph $G^{\prime}$ in $\mathcal{L}^{G}_{t}$, using Eq.(47).
@@
    If for all satellite $i$ we have $\gamma_{\text{test}}^{i}<1$, we accept the null hypothesis, and otherwise we reject the null hypothesis. We identify the fault satellite $s_{f}$ as the one with the smallest $\gamma_{\text{test}}^{i}$.
@@
Note that the above test only considers a maximum of one fault satellite. One possible extension to the test to consider multiple fault satellites is to run the algorithm iteratively, removing the detected fault satellite from the graph after each iteration until no fault is detected. Alternatively, we can consider ${}_{N_{s}}C_{N_{f}}$ combinations of fault satellites, where $N_{f}$ is the number of fault satellites, and compute the test for each combination. The detection performance in the case of multiple faults will be performed in future work.

### v3

@@ 4.1.1 Definitions
Based on the graph and observed ranges $r_{ij}$, we define a weighted graph $G=\langle V,E,W\rangle$. A realization (embedding) of $G$ is $d$-dimensional Euclidean distance space, $\mathbb{R}^{d}$, is defined as a function $\Phi$ that maps $V$ into $\mathbb{R}^{d}$, such that for each edge $e=(v_{i},v_{j})\in E$, $W(e)=r_{ij}$ [15]. A weighted graph $G=\langle V,E,W\rangle$ is called d-embeddable when there is a realization (embedding) $\Phi:V\rightarrow\mathbb{R}^{d}$ that maps the edges to points ($\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$) in $\mathbb{R}^{d}$ space.
We define a “fault satellite” as a satellite that experiences a clock phase jump, and corrupts at least one range measurement containing itself. A weighted graph is fault disprovable if and only if the embeddability of $G$ implies that it contains no fault satellites, and the non-embeddability of $G$ implies that it contains at least one fault satellite.
A graph is called rigid if it has no continuous deformation other than rotation, translation, and reflection while preserving the relative distance constraints between the vertices [20]. Otherwise, it is called a flexible graph. A graph is called k-vertex redundantly rigid if it remains rigid after removing any (k1)(k-1) vertices.
@@ 4.1.2 Embeddability of Graphs with Fault Satellites
Suppose there is at most k1k-1 fault satellites on a graph $G=\langle V,E,W\rangle$ that is $k$-vertex ($k\geq 2$) redundantly rigid in $\mathbb{R}^{d}$. If $G$ is $d$-embeddable, then $G$ contains no fault vertex (satellite) with probability 1.
@@
Given a weighted graph $G$, if $G$ is $k$-vertex redundantly rigid ($k\geq 2)$, then $G$ is fault disprovable at probability 1, if there are at most k1k-1 fault satellites.
@@
For an arbitrary graph $G$, if $G$ is not embeddable, then $G$ contains at least one fault satellite. Therefore, with Proposition 4.1, if $G$ is $k$-vertex redundantly rigid ($k\geq 2)$, then $G$ is fault disprovable at probability 1, if there are at most k1k-1 fault satellites. ∎
@@ 4.2 Properties of the Euclidean Distance Matrix
In the previous section, we proved that by using vertex redundantly rigid graphs, we can detect faults by analyzing their embeddability. Analyzing the embeddability of an arbitrary graph is known to be NP-hard [9], but if the graph is fully connected, we can evaluate it by analyzing the ranks of the GCEDM. Note that fully connected graphs of more than 5 nodes are 2-vertex redundantly rigid in $\mathbb{R}^{3}$.
@@ 4.2.1 Geometric Centered Euclidean Distance Matrix
where $\mathbf{1}\in\mathbb{R}^{n}$ is the ones vector. When no fault or noise is present $(m=0)$, the $\mathbf{G}^{n,d,0}$ will be positive semi-definite, and its rank will satisfy $\mathrm{rank}(\mathbf{G}^{n,d,0})=\mathrm{rank}(\mathbf{X}^{\top}\mathbf{X})\leq d$ [7].
@@ 4.2.2 Rank of the Euclidean Distance Matrices with Fault Satellites
[18] observed that there are more than three nonzero singular values when the GCEDM is constructed from an EDM with fault satellites. An example is provided in Figure 4. However, mathematical proofs were not provided in their paper. In this section, we provide and prove several properties of the EDM and GCEDM that are corrupted with faults and measurement noises.
@@
Figure 4: The log plot of the singular values of an example geometric centered EDM for $n=12,d=3$. The singular values are ordered in descending order of their magnitude. The blue circles show the singular values when there is no fault or noise ($\mathbf{G}^{12,3,0}$), the orange squares show the singular values when there is noise (of scale $\sigma_{w}=10^{-6}$) added to the range measurements ($\tilde{\mathbf{G}}^{12,3,0}$), and the green crosses show the singular values when both noise and fault (of scale $\bar{b}=10^{-3}$) is added ($\tilde{\mathbf{G}}^{12,3,2}$). The 12 points are identical for all cases, which are sampled randomly inside a unit cube. When noise is present, the singular values increase except for the last singular value, which is close to zero (theoretically zero). When $m$ faults with magnitudes that are larger than the noise magnitude are added, the singular values up to the $d+2m$ index increase compared to the non-fault case.
@@ 4.2.3 Distribution of the 4th Singular Value of the Geometric Centered EDM
where $\lambda_{i}$ is the $i$th singular value of the geometric centered EDM. If the ISR measurements are completely noiseless, we can detect if a fault satellite exists within the graph by checking if $\gamma_{\text{test}}$ of the GCEDMs is not 0, since GCEDMs have rank 3 when no fault exists. However, when the ISR measurements are noisy, the 4th singular value increases regardless of faults, as shown in Figure 4 and Corollary 4.4.1. Therefore, to detect faults in the presence of noise, we need to set a threshold based on the singular value distributions of the noisy (but non-fault) GCEDMs to separate the fault from the noise.
For $n=5$, we are able to compute the distribution of the test statistic analytically. The distribution of the test statistic $\gamma_{\text{test}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}$ follows a centralized chi-squared distribution with degree of freedom 1, as follows:
$\frac{\gamma_{\text{test}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}}$ $\sim\chi^{2}\left(1,0\right)\quad\text{for }n=5$ (46)
where the squared scale of the test statistic $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}$} is given by the following equation:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{a=1}^{2}\sum_{b=1}^{2}\sum_{i=1}^{5}\sum_{j=1}^{5}\left(\sigma_{m}\mathbf{\tilde{D}}_{ij}\right)^{2}\cdot\left(\hat{U}_{i,a}\hat{V}_{j,b}+\hat{U}_{j,a}\hat{V}_{i,b}\right)^{2}}$ (47)
@@
Above, $\tilde{\mathbf{G}},\tilde{\mathbf{D}}\in\mathbb{R}^{5times5}$ are the (noisy) GCEDM and the distance matrix constructed from the measured ranges, respectively. The derivation is given in Appendix F. The squared scale of the distribution is proportional to the noise variance $\sigma_{m}^{2}$ (Figure 5). Equation (47) can be modified to compute $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}$} when the variance of noise is different for each measurement, by replacing $\sigma_{m}$ with $(\sigma_{m})_{ij}$ (the noise for each range measurement).
When there is a clock jump at satellite $s_{f}$, the distribution of the test statistic $({\gamma_{\text{test}}})_{s_{f}}$ follows a non-central chi-squared distribution with 1 degree of freedom, with its squared scale $s^{2}$ given by the same equation as above, and non-centrality parameter $\lambda_{s_{f}}$:
$\frac{({\gamma_{\text{test}}})_{s_{f}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}}$ $\sim\chi^{2}\left(1,\lambda_{s_{f}}\right)\quad\text{for }n=5$ (51)
@@
$\lambda_{s_{f}}$ $=\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}}\|\hat{\mathbf{U}}^{\top}\left(\mathbf{F}\circ\mathbf{D}\right)\hat{\mathbf{V}}\|_{F}^{2}$ (52)
$\mathbf{F}$ $=\begin{cases}f_{ij}&i=s_{f}\ \text{or}\ j=s_{f}\\ 0&\text{otherwise}\end{cases}$ (53)
where $\|\cdot\|_{F}$ represents the Frobenius norm. This indicates that when all biases added by faults have the same magnitude, the non-centrality parameter $\lambda$ is proportional to the square of the fault magnitude as shown in Figure 6 and Figure 7. The derivation is given in Appendix G.
@@
Figure 5: The distribution of the test statistic $\gamma_{\text{test}}$ and its scale $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}$} (squared scale of $\lambda_{4}$) for different noise levels $\sigma_{m}$. The distribution of the test statistic $\gamma_{\text{test}}$ follows a chi-squared distribution with 1 degree of freedom. The scale $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}$} is proportional to the noise magnitude $\sigma_{m}^{2}$.
@@
$MDB(i)=\sqrt{\frac{\bar{\lambda}}{\lambda_{i}}\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}}\quad\text{where}\quad P[\chi^{2}(1,\bar{\lambda})\leq\chi^{2}_{\alpha}(1,0)]=1-\gamma$ (54)
where $\alpha$ is the false alarm rate, $\gamma$ is the detection power, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}$} is the squared scale provided in (47), and $\lambda_{i}$ is the non-centrality parameter for satellite $i$ fault case, provide in equation (52).
Figure 8 shows the computed minimum detectable biases for the two baseline methods (ephemeris comparison and data snooping) and the EDM-based fault detection, using the same geometry as Figures 5 and 6. The MDB of the data snooping marks the lowest due to its optimality, while the MDB of the ephemeris comparison method largely depends on the ephemeris accuracy. The MDB of the EDM-based detection depends on the relative geometry of each point with respect to other points, which affects the non-centrality parameter $\lambda_{sf}$.
Figure 8: The comparison of minimum detectable biases for different fault detection methods. The topology of the points and links is the same as Figure 5. Detection power $\gamma$ is fixed to 0.8, and the noise level is fixed to 1e-8. $\lambda_{s}$ in the figure reprensents $\lambda_{sf}$ in Equation 53.
If desired, the test statistic can be extended for $n\geq 6$ as:
$\lambda_{test}$ $=\sum_{i=4}^{n-1}\lambda_{i}^{2}$ (55)
The readers are referred to Appendix H for the distribution of this test statistic under nominal conditions and the existence of faults.
@@ 4.3.1 Clique Listing
In order to construct an EDM and GCEDM, we need a set of range measurements between all pairs of $n$ satellites. However, due to the occultation by the planetary body and attitude constraints, the satellites in the constellation are not fully connected with ISRs in most cases. Therefore, the first step of the fault detection algorithm is to find a set of $k$-clique subgraphs, which are fully connected subgraphs of $k$ satellites. For the graph to be 2-vertex redundantly rigid, and to use the derived distribution (46), we are interested in cliques with $k=5$. An example of 5-cliques subgraphs is shown in Figure 9.
Various exact algorithms to find all $k$-cliques have been proposed [22]. In this paper, we used the Chiba-Nishizeki Algorithm (Albo) due to its simplicity of implementation [5]. For each node $v$, Arbo expands the list of $k$-cliques by recursively creating a subgraph induced by $v$’s neighbors. The readers are referred to [22] for the pseudocode and summary of the $k$-clique listing algorithm. Note that the clique finding algorithm does not have to be executed online. Since the availability of the ISRs can be predicted beforehand when scheduling the ISLs using a coarse predicted orbit, we can estimate the topology of the ISLs for future time epochs. We can compute the list of $k$-cliques of the predicted topologies for future time epochs (e.g., one orbit), and uplink them to satellites intermittently. If some of these links were actually not available in orbit for some reason, we can remove the $k$-cliques that contain the missing links from the list.
@@ 4.3.2 Online Fault Detection
$H_{i}:$ $\text{The satellite }i\text{ is a fault satellite}$
@@
    First, we compute the scaled test statistic $\gamma_{test,G^{\prime}}=\lambda_{4}^{2}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}$ for each 5-clique subgraph $G^{\prime}$ in $\mathcal{L}^{G}_{t}$, using Eq.(47).
@@
    If for each satellite $i$ that has $\gamma_{\text{test}}^{i}<1$, we accept the null hypothesis, and otherwise we reject the null hypothesis. We identify the fault satellite $s_{f}$ as the one with the smallest $\gamma_{\text{test}}^{i}$.
@@
The above test only considers a maximum of one fault satellite. One possible extension to the test to consider multiple fault satellites is to run the algorithm iteratively, removing the detected fault satellite from the graph after each iteration until no fault is detected [19]. Alternatively, we can consider $_{N_{s}}C_{N_{f}}$ combinations of fault satellites, where $N_{f}$ is the number of fault satellites, and compute the test for each combination. The detection performance in the case of multiple faults will be performed in future work.
One practical consideration of the proposed approach is that, for satellite constellations, the fourth singular value of the GCEDM can become extremely small ($\sim 10^{-7}$) due to the large inter-satellite distances (in orders of $10^{6}$ m) compared to range measurement accuracy (in orders of sub-meters to meters). This can lead to values that approach machine precision when operating in single precision, necessitating the use of double-precision arithmetic.
4.3.3 Fault Detectable Graphs with Ephemeris Augmentation
In scenarios where the links are sparse or the number of satellites is small, the graph may lack a sufficient number of 5-cliques to perform robust fault detection using only ISR measurements. To mitigate this, we can augment the graph by “filling in” missing edges using range estimates derived from satellite ephemerides. This hybrid approach utilizes a combination of measured edges (ISR) and computed edges (ephemeris) to construct redundantly rigid subgraphs for fault detection.
However, a fault (specifically a clock phase jump) appears only in the physical ISR measurements and does not affect the ranges computed purely from ephemeris data. Consequently, for an augmented graph to reliably detect faults, the topology of the measured edges must satisfy a specific connectivity condition.
Proposition 4.5.
Consider a redundantly rigid weighted graph $G=\langle V,E,W\rangle$, where the edge set is partitioned into measured edges $E_{m}$ (ISR measurements) and computed edges $E_{e}$ (derived from ephemeris), such that $E=E_{m}\cup E_{e}$. Let a fault at satellite $v\in V$ be defined as a bias that corrupts the weights of all incident edges in $E_{m}$. The graph $G$ is fault disprovable (i.e., the graph becomes unrealizable under a fault) if and only if every vertex $v\in V$ is incident to at least one measured edge in $E_{m}$.
Proof.
See Appendix I. ∎
We denote the $k$-node subgraphs that satisfy the connectivity condition in Proposition 4.2 (i.e., every vertex is incident to at least one measured edge) as fault-detectable $k$-node subgraphs.
In scenarios where the graph is sparse or the number of satellites is small, the standard 5-clique list $\mathcal{L}^{G}_{t}$ described in Section 4.3.2 may be empty or insufficient for robust detection. In such cases, we replace $\mathcal{L}^{G}_{t}$ with the list of fault-detectable 5-node subgraphs, denoted as $\mathcal{L}^{FD}_{t}$. These subgraphs are constructed by identifying sets of 5 nodes and “filling in” missing measured edges with computed edges derived from the ephemeris to form fully connected 5-cliques.
To account for the difference in error characteristics between measured and computed edges, we modify the computation of the test statistic variance $s^{2}$. In Equation (50) (and the generalized form in Equation (117)), the uniform measurement noise variance $(\sigma_{m})_{ij}^{2}$ is replaced by an edge-dependent variance $(\sigma_{link})_{ij}^{2}$:
$(\sigma_{link})_{ij}^{2}=\begin{cases}\sigma_{m}^{2}&\text{if }(i,j)\in E_{m}\quad\text{(Measured Edge)}\\ 2\sigma_{r}^{2}&\text{if }(i,j)\in E_{e}\quad\text{(Computed Edge)}\end{cases}$ (61)
where $\sigma_{m}^{2}$ is the variance of the ISR measurement noise and $\sigma_{r}^{2}$ is the variance of the satellite position error (assuming independent position errors, the variance of the computed range error is approximately $2\sigma_{r}^{2}$).
This substitution drastically increases the number of subgraphs available for fault detection, particularly in sparse constellations or during orbital phases with limited visibility. It will also enable the detection of ephemeris faults. However, this improved availability comes at a cost: the minimum detectable bias (MDB) for these subgraphs becomes sensitive to the magnitude of ephemeris errors ($\sigma_{r}$), rather than being driven solely by the typically smaller measurement noise ($\sigma_{m}$).